# SIMPLE: Simultaneous Multi-Plane Self-Supervised Learning for Isotropic MRI Restoration from Anisotropic Data

### v1

@@ 3.2 Training
Figure 2 illustrates the training process step-by-step using the dashed red arrows. The key components of the training involve isotropic sampling, detailed in section 3.1.1, performing single-plane super-resolution on the axial and coronal planes as an initial step to generate paired data for the conditional model, and ultimately, training the generator and discriminators.
@@ 3.2.1 Single Plane Super Resolution
Although our database contains axial and coronal MRI sequences for each case, these sequences were scanned at different times. Consequently, the physiological motion of digestive tract organs (such as the bowel) and patient movement during scanning can cause shifts in the area within the field of view (FOV) of these sequences. This leads to motion artifacts and results in incomplete and inconsistent coverage of the patient anatomy between these two sequences. As a result, we cannot rely on these sequences as ground truth for our model and must address the problem using an unsupervised method. However, to still provide real samples to our model, we perform a single image deep-learning-based super-resolution model on the interpolated isotropic low-resolution MRI volume for each plane separately. This approach yields two isotropic volumes, each with high resolution in only one plane. For the single-image super-resolution model, we choose ATME, a recent approach for image-to-image translation tasks that integrates elements from GAN and diffusion models [24]. We train two ATME models, each on a different plane, and evaluate them on slice samples from the interpolated isotropic volume in the coronal and axial planes.
@@ 3.2.2 Training Setup
The generator and the two discriminators are trained together for 100 epochs with a batch size of 16. The learning rate is initially set to 0.0002 and updated using a step scheduler. The Adam optimizer is employed to update the model parameters, with beta1 set to 0.5 and beta2 set to 0.999.
@@ 4.1 Dataset
We obtained the appropriate Institutional Review Board (IRB) approval for retrospective data analysis. We collected 115 consecutive 2D MRI scans of Crohn’s Disease (CD) patients from a local hospital. All the MRI scans were acquired with a GE Medical Systems 3T MRI scanner. For the study, we utilize the coronal and axial FIESTA sequences. In both Coronal and Axial cases, the slice spacing and slice thickness are both set at 5mm, ensuring there is no overlap between adjacent slices. The average pixel spacing for the coronal sequences ranged from 0.7 mm to 0.94 mm in the x and y directions (mean 0.78 mm), while for the axial sequences, it ranged from 0.66 mm to 0.94 mm (mean 0.76 mm).
Since the number of slices varies for each case, the 3D size of the resulting isotropic volume differs accordingly. To standardize the isotropic volume to a uniform size of 512 in each dimension, we apply zero padding by adding empty slices along the edges of the coronal plane during the evaluation of our model.
Refer to captionInterpolationSMORE4SIMPLEAxial Volume
Figure 5: Fourier representation of axial slices generated from an anisotropic coronal volume. The Fourier representation of the axial slice produced by SIMPLE shows a more consistent distribution of the Fourier coefficients across the two dimensions of the image, whereas both the interpolated image (left) and the high-resolution SMORE4 image exhibit sampling artifacts. Compared to the closest slice from the axial volume (right), SIMPLE produces a similar Fourier representation, although it differs in the image domain due to the inconsistent field of view.
@@ 4.2 Comparison Methods
We compare SIMPLE to linear interpolation and the SMORE4[20] model using all performance metrics. All methods were trained and evaluated on the same coronal anisotropic MRI volumes.
@@ 4.3 Qualitative Methods
We evaluate SIMPLE’s qualitative performance through visual assessment and senior radiologist feedback. Our methods include examining generated slices in three-plane views, performing straight multi-planar reconstructions of the small intestine, analyzing the Fourier domain, and obtaining Likert scale rankings for 10 cases from our dataset.
@@
We examine sampled slices in the coronal, axial, and sagittal planes for SIMPLE and five comparison methods, evaluated on coronal or axial anisotropic volumes. Due to the lack of 3D isotropic volumes in our dataset, we do not have ground truth slices for comparison.
@@
We use an in-house application to create straight multi-planar reconstructions (MPRs) [14] of the small intestine, allowing assessment of voxel correlations in the 3D domain. We utilized a dataset of Terminal Ileum (TI) centerlines from Crohn’s Disease (CD) patients, manually annotated by an attending radiologist on the same MRI database from our local hospital.
@@
To analyze the frequency components of the generated slices, we apply the Fourier transform to sampled axial slices from the generated isotropic volume. This evaluates the preservation of high-frequency details, such as textures and edges, assessing the similarity between natural images and MRI images in the axial plane.
@@
A senior radiologist ranked slices from 10 cases in our dataset across the coronal, axial, and sagittal views. These slices were sampled from isotropic volumes generated by SIMPLE, SMORE4, and linear interpolation. The radiologist, blinded to the methods used, assigned rankings on a Likert scale from 1 to 5: 1 for non-diagnostic images, 2 for poor quality, 3 for partial visualization of anatomy with blurry borders, 4 for good visualization with mild blurring, and 5 for excellent visualization.
@@ 4.4 Quantitative Methods
We quantitatively compared our method with SMORE4 and linear interpolation using distribution-based metrics: Kernel Inception Distance (KID) and Inception Score (IS) [2], derived from the InceptionV3 network. Additionally, we evaluated the Fréchet Inception Distance (FID) [15] using the VGG16 architecture [23] as the backbone. We computed FID on overlapping tiles of 96×96 pixels with a stride of 32 pixels. To reduce inaccuracies from linear interpolation, we selected slices approximating the midpoint between two original coronal slices. For the axial FID calculation, we chose axial slices whose physical locations are closest to the original axial sequence. This ensures that the generated axial distribution is compared to slices representing the same volume in physical space.
@@ 5.1 Comparison of Ours to Baselines
Figure 3 presents views of the three primary planes of isotropic MRI volumes generated by SIMPLE and five other methods. SIMPLE effectively enhances quality and resolution in both axial and coronal planes, unlike methods that favor a specific plane. Interpolation and ATME improve the plane quality corresponding to their original anisotropic acquisition. Both SIMPLE and SMORE4 improve resolution and quality across all planes but SIMPLE excels in the axial plane by better detecting and sharpening edges and improving contrast. In the coronal plane, the differences between SIMPLE and SMORE4 are subtle, with some organ locations differing. Additionally, SMORE4 is sensitive to organ shadows and emphasizes noise. Although SIMPLE has no constraints on the sagittal plane, it improves resolution similarly to SMORE4, which does have such constraints.
@@
Figure 4 depicts depicts three straight MPRs based on four volumes: the original anisotropic volume and isotropic volumes generated by interpolation, SMORE4, and SIMPLE. SIMPLE’s reconstruction appears smoother compared to both the anisotropic and interpolated isotropic volumes, which contain discontinuities due to the larger gap between slices and inadequate generation. In some cases, the straight MPR of SMORE4 appears blurrier than that of SIMPLE.
@@
Figure 5 shows the Fourier domain of axial slices from isotropic volumes generated by interpolation, SMORE4, and SIMPLE models. SIMPLE’s Fourier transform resembles that of natural images, with a balanced mix of high and low frequencies. In contrast, the interpolated slices exhibit horizontal stripes due to the anisotropic resolution in the axial plane. Similarly, SMORE4 displays a prominent horizontal stripe near the center. All methods show distinct vertical and horizontal lines around the center, likely caused by zero-padding applied to the isotropic volume.
@@
Table 1 presents the average and standard deviation of Likert scale rankings for each main plane across different methods: interpolation, SMORE4, and SIMPLE. As expected, linear interpolation received the lowest average rank. For the coronal plane, SMORE4 and SIMPLE achieved similar ranks, with SIMPLE having a slight advantage. In the axial plane, SIMPLE’s average rank is significantly higher than SMORE4’s. In the sagittal plane, both methods received low ranks, but SIMPLE’s rank is higher.
Interpolation SMORE4 SIMPLE
Coronal Axial Coronal Axial Coronal Axial
KID 3.829 48.185 0.894 40.554 2.383 32.487
26.007 20.724 17.435
IS 0.031 0.023 0.001 0.033 0.003 0.006
0.027 0.017 0.004
FID 17.292 28.964 14.897 23.938 16.834 21.316
23.128 19.417 19.075
Table 2: Distribution-based metrics for sampled axial and coronal slices from the generated isotropic volume. It includes the scores for each plane individually, as well as the average score for both planes. Lower scores indicate better performance for all metrics.
@@
Table 2 presents the metric scores for each of the three methods. For the coronal plane, SIMPLE scores lower than interpolation but higher than SMORE4 across all metrics. In the axial plane, SIMPLE achieves the lowest scores. Overall, SIMPLE has the lowest average scores for both planes. SMORE4 performs well in the coronal plane but struggles in the axial plane, resulting in a higher score ratio between the planes compared to the more balanced ratio for SIMPLE.
@@ 5.2 Ablation Study
To evaluate the impact of individual design components in our model, we conducted two ablation studies, each testing the performance of a single discriminator. Table 3 presents the KID and IS scores for these models compared to our SIMPLE model. Model (a) achieves the lowest KID score in the coronal plane, while model (b) achieves a value similar to SIMPLE. For the IS metric, all models have similar coronal plane scores, with model (b) again close to SIMPLE. These results indicate that the axial discriminator contributes more than the coronal discriminator. This outcome was expected, as the input volume was acquired in the coronal plane, leading to a higher initial resolution in the coronal plane compared to the axial plane in the anisotropic volume. Yet, the combination of both discriminators in SIMPLE resulted in the smallest averaged KID distance reflecting the importance of the simultaneous optimization.
@@ 6 Conclusion
We introduce “SIMPLE,” a Simultaneous Multi-Plane Self-Supervised Learning approach for isotropic MRI restoration from anisotropic data. Our approach leverages anisotropic clinical data acquired in different planes, avoiding reliance on simulated downsampling. By considering the inherent three-dimensional nature of MRI data, SIMPLE generates realistic isotropic volumes, improving upon traditional through-plane slice enhancements. Our method’s flexibility allows it to be extended across various contrast types and clinical acquisition methods.
Our experimental results demonstrate that SIMPLE outperforms state-of-the-art methods both quantitatively, as measured by KID, FID, and IS, and semi-quantitatively, through evaluations by radiologists. The generated isotropic volumes facilitate more accurate volumetric analysis and 3D reconstructions, indicating significant potential for enhancing clinical diagnostic capabilities. Future work should explore further optimizations and applications of our approach to other imaging modalities and conditions.

### v2

@@ 3.2 Training
Figure 2 illustrates the training process step-by-step using the dashed arrows. The key components of the training involve isotropic sampling, detailed in section 3.1.1, performing single-plane super-resolution on the coronal, axial and sagittal planes as an initial step to generate paired data for the conditional model, and ultimately, training the generator and discriminators.
@@ 3.2.1 Single Plane Super Resolution
Although our databases contain MRI sequences in two or three planes for each case, these sequences were scanned at different times. Consequently, the physiological motion organs (such as the bowel in the digestive tract) and patient movement during scanning can cause shifts in the area within the field of view (FOV) of these sequences. This leads to motion artifacts and results in incomplete and inconsistent coverage of the patient anatomy between these two sequences. As a result, we cannot rely on these sequences as ground truth for our model and must address the problem using an unsupervised method. However, to still provide real samples to our model, we perform a single image deep-learning-based super-resolution model on the interpolated isotropic low-resolution MRI volume for each plane separately. This approach yields isotropic volumes, each with high resolution in only one plane. For the single-image super-resolution model, we choose ATME, a recent approach for image-to-image translation tasks that integrates elements from GAN and diffusion models [26]. We train two ATME models, each on a different plane, and evaluate them on slice samples from the interpolated isotropic volume in the coronal, axial and sagittal planes.
@@ 3.2.2 Training Setup
The generator and the discriminators are trained together for 100 epochs with a batch size of 16. The learning rate is initially set to 0.0002 and updated using a step scheduler. The Adam optimizer is employed to update the model parameters, with beta1 set to 0.5 and beta2 set to 0.999.
Refer to captioncase 1case 2case 3(a)(b)(c)(d)(e)
Figure 4: Straight multi-planar reconstructions (MPR) of the terminal ileum in Crohn’s disease patients for three cases based on coronal volumes: (a) anisotropic volume, (b) linearly interpolated isotropic volume, (c) averaged ATMEs isotropic volume, (d) SMORE4 isotropic volume, and (e) SIMPLE isotropic volume. The MPRs generated from the SIMPLE reconstructed volume were less prone to artifacts stemming from the anisotropic MRI acquisition.
@@ 4.1 Dataset
We utilize two distinct datasets to evaluate our model across different anatomical organs, variations in spatial resolution and volume dimensions, as well as different levels of slice-to-slice motion artifacts.
Abdomen Dataset
We obtained the appropriate Institutional Review Board (IRB) approval for retrospective data analysis. We collected 115 consecutive 2D MRI scans of Crohn’s Disease (CD) patients from a local hospital. All the MRI scans were acquired with a GE Medical Systems 3T MRI scanner. For the study, we utilize the coronal and axial FIESTA sequences. In both Coronal and Axial cases, the slice spacing and slice thickness are both set at 5mm, ensuring there is no overlap between adjacent slices. The average pixel spacing for the coronal sequences ranged from 0.7 mm to 0.94 mm in the x and y directions (mean 0.78 mm), while for the axial sequences, it ranged from 0.66 mm to 0.94 mm (mean 0.76 mm). To standardize the isotropic volume to a uniform size of 512 in each dimension, we apply zero padding by adding empty slices along the edges of the coronal plane during the evaluation of our model.
Oasis-1 Dataset
We use the publicly available OASIS-1 dataset, containing 416 cross-sectional MRI brain scans from adults across different age groups and cognitive statuses. Each subject’s data includes 3-4 T1-weighted MRI scans in the sagittal plane, with 1.25 mm slice spacing and 1 mm pixel resolution. The dataset also provides an averaged, co-registered image resampled to an isotropic voxel size of 1 mm, serving as the 3D isotropic ground truth for our study. From this, we generate three anisotropic 2D MRI image sequences in the coronal, axial, and sagittal planes, with a slice spacing of 3 mm and a pixel resolution of 1 mm, by applying a Gaussian filter ($\sigma=3$) and selecting every third slice. Zero padding is applied to standardize the size to 256 in each dimension.
Refer to captionInterpolationSMORE4Avg. ATMESIMPLEGT Volume
Figure 5: Fourier representation of axial and coronal slices generated from an anisotropic axial and coronal volumes, accordingly. SIMPLE’s Fourier representation shows a more consistent distribution of the Fourier coefficients across the two dimensions of the image, whereas other methods images exhibit sampling artifacts. Compared to the closest slice from the axial or coronal GT volume, SIMPLE produces a similar Fourier representation.
@@ 4.2 Comparison Methods
We compare SIMPLE against linear interpolation, SMORE4 [22], and an approach similar to Yucheng et al. [17], in which super-resolution 3D volumes are generated by ATME in each plane and then averaged. All methods were evaluated on identical anisotropic MRI volumes from the same plane using consistent performance metrics. For evaluation, we used the coronal plane for the abdomen dataset and the sagittal plane for the brain dataset.
@@ 4.3 Qualitative Methods
We assess SIMPLE’s qualitative performance through visual assessments, multi-planar reconstructions, Fourier domain analysis, and senior radiologists feedback.
@@
We examine sampled slices in the coronal, axial, and sagittal planes for SIMPLE and comparison methods, evaluated on anisotropic volumes. Unlike the brain dataset, the abdomen dataset lacks 3D isotropic volumes, so ground truth slices for comparison are not available.
@@
For the abdomen dataset, we use an in-house application to create straight multi-planar reconstructions (MPRs) [15] of the small intestine, allowing assessment of voxel correlations in the 3D domain. We utilized a dataset of Terminal Ileum (TI) centerlines from Crohn’s Disease (CD) patients, manually annotated by an attending radiologist on the same MRI database from our local hospital.
@@
To analyze the frequency components of the generated slices, we apply the Fourier transform to sampled slices from the isotropic volume in planes other than the evaluation plane. This evaluates the preservation of high-frequency details, such as textures and edges, assessing the similarity between ground-truth and generated MRI images.
@@
A senior radiologist ranked slices from 10 cases in our abdomen dataset across the coronal, axial, and sagittal views. These slices were sampled from isotropic volumes generated by SIMPLE, SMORE4, and linear interpolation. The radiologist, blinded to the methods used, assigned rankings on a Likert scale from 1 to 5: 1 for non-diagnostic images, 2 for poor quality, 3 for partial visualization of anatomy with blurry borders, 4 for good visualization with mild blurring, and 5 for excellent visualization.
@@ 4.4 Quantitative Methods
We quantitatively compared SIMPLE to other methods using distribution-based metrics. For the abdomen dataset we used Kernel Inception Distance (KID) and Inception Score (IS) [3], derived from the InceptionV3 network. Additionally, we evaluated the Fréchet Inception Distance (FID) [16] using the VGG16 architecture [25] as the backbone.
In their thorough review, Yi et al. [29] highlight that standard metrics like PSNR and SSIM do not effectively capture the visual quality in GAN-generated medical images, raising concerns over the limitations of traditional full-reference metrics. Consequently, we chose to assess the brain dataset using the KID metric, which does not rely exclusively on ground-truth comparisons. We averaged the KID scores across all 2D slices in each plane. Additionally, we assessed 3D volume quality by averaging KID scores across 2D slices per plane, then averaging across cases.
@@ 5.1 Comparison of Ours to Baselines
Figures  3 presents views of the three primary planes of isotropic MRI volumes generated by all methods. SIMPLE effectively enhances quality and resolution in both axial and coronal planes, unlike methods that favor a specific plane. Interpolation and ATME improve the plane quality corresponding to their original anisotropic acquisition. SIMPLE, SMORE4 and averaged ATMEs all improve resolution and quality across all planes, but SIMPLE outperforms in the planes perpendicular to the acquisition plane by better detecting and sharpening edges and improving contrast. In the acquisition plane, the differences between SIMPLE and SMORE4 are subtle, with some organ locations differing. For the abdomen dataset, SMORE4 is sensitive to organ shadows and emphasizes noise. While SIMPLE has no constraints in the sagittal plane for the abdomen dataset, it improves resolution. For the brain dataset, differences between SIMPLE to SMORE4, and averaged ATMEs are less pronounced than in the abdomen dataset, due to smaller slice spacing and lower likelihood of organ motion.
@@
Figure 4 depicts three straight MPRs based on five volumes: the original anisotropic volume and isotropic volumes generated by all methods. SIMPLE’s reconstruction appears smoother compared to both the anisotropic and interpolated isotropic volumes, which contain discontinuities due to the larger gap between slices and inadequate generation. In some cases, the straight MPRs of SMORE4 and averaged ATMEs appear blurrier than that of SIMPLE.
@@
Figure 5 shows the Fourier domain of axial and coronal slices from isotropic volumes generated by all methods. For both datasets, SIMPLE’s Fourier transform resembles that of natural images, with a balanced mix of high and low frequencies. In contrast, the interpolated axial abdominal slices exhibit horizontal stripes due to the anisotropic resolution in the axial plane. Similarly, SMORE4 and averaged ATMEs display a prominent horizontal stripe near the center. All methods show distinct vertical and horizontal lines around the center, likely caused by zero-padding applied to the isotropic volume. For the brain dataset, it is apparent that the coronal and axial slices of the competitive methods lack high-frequency components and exhibit a greater disparity from the Fourier domain of the original images.
@@
Table 1 presents the average and standard deviation of Likert scale rankings for each main plane across all methods. As expected, linear interpolation received the lowest average rank. For the coronal plane, SMORE4 and SIMPLE achieved similar ranks, with SIMPLE having a slight advantage. In the axial plane, SIMPLE’s average rank is significantly higher than SMORE4’s. In the sagittal plane, both methods received low ranks, but SIMPLE’s rank is higher.
Interpolation SMORE4 Avg. ATME SIMPLE
Coronal Axial Coronal Axial Coronal Axial Coronal Axial
KID 3.829 48.185 0.894 40.554 0.639 37.213 2.383 32.487
26.007 20.724 18.926 17.435
IS 0.031 0.023 0.001 0.033 0.005 0.022 0.003 0.006
0.027 0.017 0.013 0.004
FID 17.292 28.964 14.897 23.938 15.387 22.969 16.834 21.316
23.128 19.417 19.178 19.075
Table 2: Distribution-based metrics for sampled axial and coronal slices from the generated isotropic volume in the abdomen dataset. It includes the scores for each plane individually and the average score for both planes. Lower scores indicate better performance.
Interpolation SMORE4 Avg. ATME SIMPLE
Cor Ax Sag Cor Ax Sag Cor Ax Sag Cor Ax Sag
2D
KID
0.89 0.964 0.386 0.47 0.495 0.354 0.418 0.455 0.241 0.396 0.422 0.231
3D
KID
30.425 29.295 29.36 28.709
Table 3: KID scores for sampled axial, coronal and sagittal slices from the generated isotropic volume in the brain dataset. The table includes averaged KID scores for each plane individually and the overall average KID score across all planes for each case.
@@
Table 2 and  3 present the metric scores for each of the methods applied to the abdomen and brain datasets, respectively. In the abdomen dataset, SIMPLE scores lower than interpolation but higher than SMORE4 and averaged ATMEs in the coronal plane across all metrics. In the axial plane, SIMPLE achieves the lowest scores. SMORE4 and averaged ATME perform well in the coronal plane but struggle in the axial, resulting in a higher score ratio between the planes compared to the more balanced ratio for SIMPLE. For the brain dataset, SIMPLE produces slightly lower KID scores than SMORE4 and averaged ATMEs in both 2D and 3D evaluations, and significantly lower scores compared to interpolation. The most notable differences across all methods are observed in the coronal and axial planes.
@@ 5.2 Ablation Study
Two ablation studies were conducted on the abdomen dataset to assess the individual contributions of each discriminator. Table 4 reports KID and IS scores for these models in comparison to SIMPLE. Model (a) achieved the lowest KID score in the coronal plane, while model (b) performed similarly to SIMPLE. IS scores across all models were comparable, with model (b) closest to SIMPLE. These results indicate that the axial discriminator contributes more significantly, likely due to the coronal plane’s higher resolution in the anisotropic data. The combined use of both discriminators in SIMPLE achieved the lowest average KID, emphasizing the importance of simultaneous optimization.
@@ 6 Conclusion
We introduce SIMPLE, a self-supervised approach for isotropic MRI restoration from anisotropic data using multi-plane clinical images. SIMPLE outperforms state-of-the-art methods in quantitative metrics and radiologist evaluations, enhancing volumetric analysis and 3D reconstructions with potential clinical benefits. Future work will focus on optimizations and applications to other imaging modalities.
Supplementary Material
[Uncaptioned image]Sagittal VolumeAxial VolumeCoronal VolumeSagittal ATMEAxial ATMECoronal ATMEAvg. ATMESMORE4SIMPLEGT
Axial Coronal Sagittal
Figure 6: Multi-plane views of slices from an isotropic MRI volume, derived from a brain dataset, generated by seven competing methods and compared with the ground-truth (GT) volume. From left to right: linear interpolation on anisotropic sagittal volume, linear interpolation on anisotropic axial volume, linear interpolation on anisotropic coronal volume, ATME on anisotropic sagittal volume, ATME on anisotropic axial volume, ATME on anisotropic coronal volume, averaged ATME of coronal, axial and sagittal planes on anisotropic sagittal volume, SMORE4 on anisotropic sagittal volume, and SIMPLE on anisotropic sagittal volume. Average ATME, SMORE4, and SIMPLE effectively enhance image quality and resolution across all planes, whereas linear interpolation and ATME demonstrate biases toward their original anisotropic acquisition plane. SIMPLE shows slightly better performance than Average ATME and SMORE4, achieving sharper images with less noise, fewer shadows, and enhanced contrast.